A 456 mL sample of argon gas is expanded at constant at constant pressure to 1.75 L at 55ºC. What was the original temperature of the gas in º C?

a.
-188 ºC

b.
188 º C

c.
200 F

d.
-200 F
1
Expert's answer
2020-08-02T15:51:32-0400

V1 / T1 = V2 / T2 ........


0.456 L / T1 = 1.75 L / 328 K.


T1 = 0.456 × 328 / 1.75 .


T1 = 85.467 K .